OpenStudy (anonymous):

a box contains 5 red balls,3 green balls and 4 blue balls. if two balls are drawn in succession without replacement.what is the probability that both are red?..help plzz..

OpenStudy (anonymous):

5/12x4/11

Directrix (directrix):

There are 5 + 3 + 4 = 12 balls in the box. Of these, 5 are red. Probability is the number of desirable outcomes divided by the number of possible outcomes. So, the probability of getting a red ball on the first draw is 5/12.

OpenStudy (anonymous):

5/33 should be the answer

Directrix (directrix):

The drawing is done without replacement. Assuming that a red ball was selected on the first draw, then there are now 4 red balls in the box. Also, there are now 11 total balls in the box. The probability of getting a red ball on the second draw is 4/11. The event for which we want the probability is the drawing of two red balls in succession. Therefore, that probability is 5/12 * 4/11 = 20/132 = 5/33. --> If you conduct this drawing of two red balls 33 different times, you can expect to get Red-Red, on average, 5 of the 33 times.

OpenStudy (anonymous):

thanks..

Directrix (directrix):

We were happy to help.